Morgan has an 81/2-foot board that must be cut into boards that are 5/6 feet long. How many of the smaller boards can he get from the 81/2-foot board?
Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

I assume that 8 1/2 is what you meant to write not 81 over 2
divide 81.5 by 5/6 which is the same as multiplying 81.5 by 6/5
example
8 divided by 1/2= 8*2=16
Think about it.
1/2 goes into 1 two times
1/2 goes into 8 sixteen times
so we have 81.5*6/5=81.5*1.2

Answer by solver91311(24713)



8 is 16/2, so 8 and 1/2 is 17/2. 17/2 is 51/6. How many times, excluding any fractional part does 5 go into 51?


John